• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 01:37:29 on December 04, 2009

    Entry number 304521

    This entry is a followup to: 304198

    keyword=cavity mon update

    After yesterday's gain adjustment on cavities, all the channels
    are not saturating at 100 uA, but one channel (roc26, slot1, chan 2)
    is VERY close to saturating -- see fig 1. The other channels were
    hovering around 220 to 250K.

    In Pan DB this channel corresponds to

    datamap     bpm   bpmcav3   adcx       8      2     bpmcav3yr
    

    We're already at a minimum gain on this one, so what we need to do
    during the break is:

    1) Reduce the electronic offset from 4 V (way too high) to 1 or 2 V.

    2) Get better at noticing this problem. It's a matter of looking at
    the Peak values of raw ADC.





    Figure 1